# Training-and-prompt-free General Painterly Harmonization Using Image-wise Attention Sharing

### v1

@@ H.1. Inpainting
We incorporate our proposed share-attention w/ reweighting into the inpainting method Repaint(Lugmayr et al., 2022). And provide the functionality of exemplar guided inpainting. We replaced the noisy latent inside mask area of $z^{\text{comp}}_{T}$ to the noise of $z^{\text{b}}_{T}$. The result can be found in Fig. 14, showing that our proposed method is able to provide current inpaint method additional exemplar information based on existing framework.
@@
Figure 14. The left column is the given $I^{\text{comp}}$ and corresponding inpaint mask, the middle column is the additionally provided $I^{\text{f}}$ (in red box) and $I^{\text{b}}$. The last column is the output image.
H.2. Semantic mixing
@@
We test the compatibility of share-attention layer with semantic mixing method InjectFusion (Jeong et al., 2024), which is the adaption of the renowned Asyrp(Kwon et al., 2022) approach. The core concept shared by these methods involves blending the semantic information from two images by manipulating their h-space, specifically the intermediary attention layer within the diffusion UNet architecture. To integrate the shared-attention layer into the InjectFusion, we simply allow the output image to attend to additionally provided $I^{\text{b}}$ via our share-attention layer. The generated result can be found in Fig. 15
@@
Figure 15. The columns, from left to right, represent $I^{\text{f}}$, $I^{\text{b}}$, and $I^{\text{o}}$, where our share-attention layer is able to perform astonishing semantic mixing when combined with corresponding method.
Figure 16. Qualitative comparison of WikiArt combined with COCO
Table 4. Quantitative comparison of WikiArt combined with COCO
Painterly Harmonization Style Transfer Harmonization
Composite Ours (low/high) TF-ICON PHDifff PHDNet SDEdit DIB QuantArt VCT StyTr2
$Venue$ ICCV’23 MM’23 AAAI’23 ICLR’22 WACV’20 CVPR’23 ICCV’23 CVPR’22
$LPIPS_{bg}\downarrow$ 0 0.08/0.10 0.21/0.34 0.08/0.11 0.34 0.36 0.11 0 0 0
$LPIPS_{fg}\downarrow$ 0 0.10/0.27 0.30/0.35 0.11/0.30 0.32 0.29 0.23 0.28 0.12 0.16
$CLIP_{img}\uparrow$ 99.99 92.88/80.40 83.79/82.65 91.78/80.32 81.65 80.70 88.43 81.22 90.86 87.43
$CLIP_{style}\uparrow$ 45.60 47.96/55.25 49.51/50.02 46.17/53.95 50.64 50.46 48.59 50.02 47.85 50.31
$CLIP_{dir}\uparrow$ 0 5.44/19.20 7.63/11.18 6.49/18.14 15.80 13.46 8.67 15.23 6.26 11.16
@@
Figure 17. More painterly harmonization result of our proposed TF-GPH on GPH Benchmark. The row 1, is the input foreground objects ,and the row 2,3,4 are the corresponding outputs, The row 5 is the input background objects , the row 6 is the composite image with given face to paste, and the row 7 is the corresponding outputs
@@
Figure 18. More style transfer result of our proposed TF-GPH on GPH Benchmark. The images inside blue box, serves as both the content reference for the corresponding row and the style reference for the corresponding column,

### v2

@@ Appendix I Appendix: More Applications
We incorporate our proposed share-attention w/ reweighting into the inpainting method Repaint(Lugmayr et al. 2022). And provide the functionality of exemplar guided inpainting. We replaced the noisy latent inside mask area of $z^{\text{comp}}_{T}$ to the noise of $z^{\text{b}}_{T}$. The result can be found in Fig. 14, showing that our proposed method is able to provide current inpaint method additional exemplar information based on existing framework.
@@
Figure 14: The left column is the given $I^{\text{comp}}$ and corresponding inpaint mask, the middle column is the additionally provided $I^{\text{f}}$ (in red box) and $I^{\text{b}}$. The last column is the output image.
Semantic mixing
@@
We test the compatibility of share-attention layer with semantic mixing method InjectFusion (Jeong, Kwon, and Uh 2024), which is the adaption of the renowned Asyrp(Kwon, Jeong, and Uh 2022) approach. The core concept shared by these methods involves blending the semantic information from two images by manipulating their h-space, specifically the intermediary attention layer within the diffusion UNet architecture. To integrate the shared-attention layer into the InjectFusion, we simply allow the output image to attend to additionally provided $I^{\text{b}}$ via our share-attention layer. The generated result can be found in Fig. 15
@@
Figure 15: The columns, from left to right, represent $I^{\text{f}}$, $I^{\text{b}}$, and $I^{\text{o}}$, where our share-attention layer is able to perform astonishing semantic mixing when combined with corresponding method.
Figure 16: Qualitative comparison of WikiArt combined with COCO
Table 6: Quantitative results of GPH-Benchmark ( represents the method with inference-time-adjustable hyperparameters. The left side of / represents content emphasis strategy, while the right side of / represents stylized emphasis strategy.)
Painterly Harmonization (256x256) Style Transfer (256x256)
Ours ArtoPIH ProPIH TF-ICON PHDiff Ours StyleID Z-STAR StyTr2
$Venue$ AAAI’24 AAAI’24 ICCV’23 MM’23 - CVPR’24 CVPR’24 CVPR’22
$LP_{\text{bg}}\downarrow$ 0.07/0.07 0.05 0.06/0.06 0.16/0.32 0.06/0.06 0.66/0.47 0.55/0.56 0.63/0.63 0.55
$LP_{\text{fg}}\downarrow$ 0.04/0.26 0.22 0.18/0.30 0.29/0.33 0.06/0.31 0.05/0.40 0.29/0.42 0.12/0.33 0.33
$CP_{\text{img}}\uparrow$ 97.58/84.41 90.58 93.30/83.81 90.26/87.85 97.46/73.65 98.06/75.55 84.65/71.64 93.33/71.18 84.32
$CP_{\text{style}}\uparrow$ 48.74/54.65 50.82 49.53/52.43 48.46/48.48 49.04/53.96 61.38/75.52 69.22/78.96 59.61/69.50 64.98
$CP_{\text{dir}}\uparrow$ 0.01/9.11 3.69 2.03/10.09 2.66/3.91 0.35/12.39 2.08/45.02 24.92/47.68 8.44/32.84 20.42
Table 7: Quantitative results of WikiArt w/ COCO ( represents the method with inference-time-adjustable hyperparameters. The left side of / represents content emphasis strategy, while the right side of / represents stylized emphasis strategy.)
Painterly Harmonization (512x512)
Ours ArtoPIH ProPIH TF-ICON PHDiff PHDNet SDEdit DIB
$Venue$ AAAI’24 AAAI’24 ICCV’23 MM’23 AAAI’23 ICLR’22 WACV’20
$LP_{\text{bg}}\downarrow$ 0.08/0.10 0.24 0.29/0.29 0.21/0.34 0.08/0.11 0.34 0.36 0.11
$LP_{\text{fg}}\downarrow$ 0.10/0.27 0.30 0.25/0.37 0.11/0.30 0.10/0.39 0.32 0.29 0.23
$CP_{\text{img}}\uparrow$ 92.88/80.40 83.55 86.94/78.33 83.79/82.65 91.78/80.32 81.65 80.70 88.4
$CP_{\text{style}}\uparrow$ 47.96/55.25 49.92 47.74/51.07 49.51/50.02 46.17/53.95 50.64 50.46 48.59
$CP_{\text{dir}}\uparrow$ 5.44/19.20 13.04 9.20/18.11 7.63/11.18 6.49/18.14 15.80 13.46 8.67
Table 8: Quantitative results of WikiArt w/ COCO ( represents the method with inference-time-adjustable hyperparameters. The left side of / represents content emphasis strategy, while the right side of / represents stylized emphasis strategy.)
Painterly Harmonization (256x256)
Ours ArtoPIH ProPIH TF-ICON PHDiff PHDNet SDEdit DIB
$Venue$ AAAI’24 AAAI’24 ICCV’23 MM’23 AAAI’23 ICLR’22 WACV’20
$LP_{\text{bg}}\downarrow$ 0.05/0.06 0.04 0.05/0.05 0.18/0.31 0.06/0.06 0.12 0.14 0.8
$LP_{\text{fg}}\downarrow$ 0.06/0.23 0.23 0.19/0.34 0.11/0.30 0.06/0.28 0.26 0.21 0.18
$CP_{\text{img}}\uparrow$ 95.83/83.31 87.31 90.84/81.62 83.79/82.65 94.28/78.66 79.64 81.33 87.63
$CP_{\text{style}}\uparrow$ 48.26/54.41 51.20 49.07/52.38 49.51/50.02 46.17/53.95 52.13 51.71 51.43
$CP_{\text{dir}}\uparrow$ 2.92/13.11 10.23 6.73/12.91 7.63/11.18 7.49/21.11 16.77 14.52 9.11
@@
Figure 17: More painterly harmonization result of our proposed TF-GPH on GPH Benchmark. The row 1, is the input foreground objects ,and the row 2,3,4 are the corresponding outputs, The row 5 is the input background objects , the row 6 is the composite image with given face to paste, and the row 7 is the corresponding outputs
@@
Figure 18: More style transfer result of our proposed TF-GPH on GPH Benchmark. The images inside blue box, serves as both the content reference for the corresponding ROW and the style reference for the corresponding COLUMN,